# Universal Performance Gap of Neural Quantum States Applied to the Hofstadter-Bose-Hubbard Model

### v2

@@ Abstract
Neural Quantum States (NQS) have demonstrated significant potential in approximating ground states of many-body quantum systems, though their performance can be inconsistent across different models. This study investigates the performance of NQS in approximating the ground state of the Hofstadter-Bose-Hubbard (HBH) model, a boson system on a two-dimensional square lattice with a perpendicular magnetic field. Our results indicate that increasing magnetic flux leads to a substantial increase in energy error, up to three orders of magnitude. Importantly, this decline in NQS performance is consistent across different optimization methods, neural network architectures, and physical model parameters, suggesting a significant challenge intrinsic to the model. Despite investigating potential causes such as wave function phase structure, quantum entanglement, fractional quantum Hall effect, and the variational loss landscape, the precise reasons for this degradation remain elusive. The HBH model thus proves to be an effective testing ground for exploring the capabilities and limitations of NQS. Our study highlights the need for advanced theoretical frameworks to better understand the expressive power of NQS which would allow a systematic development of methods that could potentially overcome these challenges.
@@ 2 Hofstadter-Bose-Hubbard model
The (noninteracting) Hofstadter model [20, 21] was introduced to describe a two-dimensional charged particle moving on a lattice pierced by a uniform perpendicular magnetic field. The interplay of the lattice periodicity with the length scale set by the magnetic field leads to the emergence of a fractal energy spectrum, as two structures – that of the Bloch bands and of the Landau levels – need to be imposed simultaneously. The model supports a topological band structure and a fractal phase diagram for the integer quantum Hall effect [22, 23, 24, 25, 26]. Its controllable physical realization is accessible in ultracold atomic experimental setups [27, 28, 29]. Thanks to a successful inclusion of particle interactions in an experiment [30], the model was upgraded to the many-body setting, paving the way towards the study of the fractional quantum Hall regime [31, 32, 33, 34].
@@ 3.3.2 Supervised imaginary time evolution
The target wave function is held fixed for multiple steps and is updated after energy decreases below a certain threshold value. The time step $\Delta\tau$ is chosen to minimize the average energy of the target wave function. A key advantage of this method is that it does not require the computation of the quantum geometric tensor, as the geometric structure of the quantum state space is inherently included within the loss function. Throughout this paper, we refer to the latter method as supervised imaginary time evolution (SITE).
@@ 3.3.3 Supervised learning of exact solution
The third method involves supervised learning directly applied to the ground state wave function $\psi_{\mathrm{ED}}(s)$ obtained through exact diagonalization (ED). This is achieved by minimizing the loss as specified in Eq. 9, but with $\psi_{\mathrm{ED}}(s)$ replacing $\psi_{\mathrm{T}}(s)$. This approach is primarily useful for testing and evaluating the performance of NQS, as having the ED solution available in practice would render the NQS solution unnecessary. Learning the wave function directly in this manner is generally insightful as it assesses whether the expressivity of neural networks is adequate for representing the relevant wave function. This approach is particularly valuable when investigating the SITE method, as it is expected to provide an upper bound for the precision of the ground state approximation. Ideally, performing imaginary time evolution as described in Eq. 8 should eventually yield the target wave function $\psi_{\mathrm{T}}(s)=\psi_{\mathrm{ED}}(s)$, resulting in the same loss function as used here. We maintained the same optimization hyperparameters as those used in the SITE method.
@@ 4.1 Drop of NQS performance in strong magnetic flux regime
We begin by examining the dependence of NQS performance on the strength of magnetic flux, utilizing SR and SITE for optimization. For this we conduct experiments with a $4\times 5$ lattice containing 4 particles and employ MLP architecture with two hidden layers, each comprising 32 neurons. This neural network has 1856 variational parameters, which is more than two times fewer than the dimension of the Hilbert space for this case, which amounts to 4845. We observed a rapid decline in the energy accuracy of NQS with increasing values of $\alpha$. Figure 1 illustrates the relationship between energy error (obtained by comparing with exact diagonalization) and $\alpha$. The energy error associated with NQS escalates by several orders of magnitude within the $0.20<\alpha<$ 0.2 range and experiences a further increase of an order of magnitude within the $0.30.2<\alpha<$ 0.3 range. Importantly, the observed decline in performance is virtually identical whether SR or SITE is used for NQS optimization. This uniformity implies that the underlying issue is unlikely to originate from the optimization algorithm itself.
@@
We also incorporated into Figure 1 the results obtained using the density matrix renormalization group (DMRG) method for comparison. For this we utilised TeNPy [58] Python package. We note that this is not the best result that could be achieved with DMRG as the virtual bond dimension of the Matrix Product State (MPS) was set to 8 to closely align the number of variational parameters (2,336) with that of our NQS. Notably, NQS shows significantly better performance at lower $\alpha$ values, becoming comparable to DMRG at higher values. While the performance of DMRG also deteriorates with increasing $\alpha$, this decline is considerably less pronounced than in the case of NQS.
We also observe a similar decrease in accuracy with increasing magnetic flux when directly maximizing the overlap between the NQS and the wave function obtained from ED (described in sec. 3.3.3). In this context, we assess performance by measuring the overlap deviation from 1, defined as:
@@
The overlap deviation from the target wave function vector increases by orders of magnitude in the regime of strong magnetic flux, compared to the scenario without a magnetic field. Remarkably, the specifics of the neural network architecture, as well as the size and density of the physical model, have minimal impact on this result. To validate this observation, we undertook an extensive series of experiments across a variety of neural network architectures, lattice dimensions, and particle counts. We explored every combination of the following parameters: the number of layers (2, 3, or 4); the number of neurons per hidden layer (32, 64, or 128); the number of lattice columns (4, 6, or 8); and the number of particles (4, 6, or 8). Additionally, we conducted analogous experiments using a CNN architecture, where we varied the number of channels (10, 20, or 44) instead of the number of neurons. The results from these experiments are illustrated in Fig. 1. Clearly, a relationship exists between the overlap deviation and the ratio of NQS parameters to the number of dimensions in the Hilbert space. The relationship is characterized by a significant dispersion of approximately one order of magnitude. This variability arises because different neural network configurations may vary in their optimality for this problem, and different physical system configurations may present varying levels of difficulty for approximation. Despite this large variation, a consistent overlap deviation gap of three to four orders of magnitude is observed between cases with $\alpha=0$ and $\alpha=0.3$.
@@
Figure 1: Left: Dependence of NQS energy error on magnetic flux. Right: Overlap deviation plotted against the ratio of NQS parameters to the dimension of the Hilbert space. Each point represents a different run, varying in neural network architecture, lattice size, and particle number. Blue and red symbols indicate runs without magnetic fields and with strong magnetic fields, respectively. Circles and crosses denote runs using MLP and CNN neural network architectures, respectively.
@@ 4.2.4 Fractional quantum Hall effect
As it is well known [22], the noninteracting Hofstadter model is a paradigmatic model featuring the emergence of topological single-particle energy bands. In the presence of strong particle interactions and in the regime of a partially filled lowest Landau band, the Hofstadter Bose-Hubbard model has been conjectured to support the fractional Chern insulating (FCI) states, analogous to the Laughlin states in continuous systems [33, 34, 59, 60]. For bosons, the principal FCI state is ecpected at the filling factor of $\nu=1/2$ particles per flux quantum. Note that in finite systems with open boundaries, the number of plaquettes pierced by the flux is $(L_{x}-1)\times(L_{y}-1)$, thus the filling factor is defined [61] as $\nu=N/[\alpha(L_{x}-1)(L_{y}-1)]$, with $N$ denoting the number of particles. The formation of a fractionalized state on matter would certainly pose an additional (and unexplored) challenge for the NQS approach. However, we argue that in our case the performance drop cannot be ascribed to the stabilization of FCI states. In Figure 6 we plot the overlap deviation on the magnetic flux $\alpha$ (left panel) and the corresponding dependence on the filling factor $\nu$ (right panel) for a selection of lattice geometries. We see that the characteristic exponential growth of the overlap deviation is mainly associated with the regime of filling factors well above unity. The behavior of the overlap deviation below $\nu=1$ is rather erratic and the emerging peaks do not consistently point to a particular value of $\nu$, as one would expect looking for signals of FCI stabilization. All in all, we believe that in our experiments the considered regimes are not favourable for the emergence of FCI states, and they cannot be identified as a source of the performance gap.

### v3

@@ Abstract
Neural Quantum States (NQS) have demonstrated significant potential in approximating ground states of many-body quantum systems, though their performance can be inconsistent across different models. This study investigates the performance of NQS in approximating the ground state of the Hofstadter-Bose-Hubbard (HBH) model, an interacting boson system on a two-dimensional square lattice with a perpendicular magnetic field. Our results indicate that increasing magnetic flux leads to a substantial increase in energy error, up to three orders of magnitude. Importantly, this decline in NQS performance is consistent across different optimization methods, neural network architectures, and physical model parameters, suggesting a significant challenge intrinsic to the model. Despite investigating potential causes such as wave function phase structure, quantum entanglement, fractional quantum Hall effect, and the variational loss landscape, the precise reasons for this degradation remain elusive. The HBH model thus proves to be an effective testing ground for exploring the capabilities and limitations of NQS. Our study highlights the need for advanced theoretical frameworks to better understand the expressive power of NQS which would allow a systematic development of methods that could potentially overcome these challenges.
@@ 2 Hofstadter-Bose-Hubbard model
The Hofstadter model [20, 21] was introduced to describe a two-dimensional charged particle moving on a lattice pierced by a uniform perpendicular magnetic field. The interplay of the lattice periodicity with the length scale set by the magnetic field leads to the emergence of a fractal energy spectrum, as two structures – that of the Bloch bands and of the Landau levels – need to be imposed simultaneously. The model supports a topological band structure and a fractal phase diagram for the integer quantum Hall effect [22, 23, 24, 25, 26]. Its controllable physical realization is accessible in ultracold atomic experimental setups [27, 28, 29]. Thanks to a successful inclusion of particle interactions in an experiment [30], the model was upgraded to the many-body setting, paving the way towards the study of the fractional quantum Hall regime [31, 32, 33, 34].
@@ 3.3.2 Supervised imaginary time evolution
The target wave function is held fixed for multiple steps and is updated after energy decreases below a certain threshold value. The time step $\Delta\tau$ is chosen to minimize the average energy of the target wave function. Throughout this paper, we refer to the latter method as supervised imaginary time evolution (SITE). Unlike SR, this method does not require computing the quantum geometric tensor, as the geometric structure of the quantum state space is inherently embedded in the loss function. Another key difference is that in SITE, the target wave function remains fixed for multiple steps, whereas in methods that directly minimize the energy, the effective target wave function changes with each step [50].
@@ 3.3.3 Supervised learning of exact solution
The third method involves supervised learning directly applied to the ground state wave function $\psi_{\mathrm{ED}}(s)$ obtained through exact diagonalization (ED). This is achieved by minimizing the loss as specified in Eq. 9, but with $\psi_{\mathrm{ED}}(s)$ replacing $\psi_{\mathrm{T}}(s)$. This approach is primarily useful for testing and evaluating the performance of NQS, as having the ED solution available in practice would render the NQS solution unnecessary. Learning the wave function directly in this manner is generally insightful as it assesses whether the expressivity of neural networks is adequate for representing the relevant wave function. This approach is particularly valuable when investigating the SITE method, as it is expected to provide an upper bound for the precision of the ground state approximation. Ideally, performing imaginary time evolution as described in Eq. 8 should eventually yield the target wave function $\psi_{\mathrm{T}}(s)=\psi_{\mathrm{ED}}(s)$, resulting in the same loss function as used here. We maintained the same optimization hyperparameters as those used in the SITE method. In the supervised learning scenario described in this section, we assess performance by measuring the overlap deviation from 1, which is defined as:
@@ 4.1 Drop of NQS performance in strong magnetic flux regime
We begin by examining the dependence of NQS performance on the strength of magnetic flux, utilizing SR and SITE for optimization. For this we conduct experiments with a $4\times 5$ lattice containing 4 particles and employ MLP architecture with two hidden layers, each comprising 32 neurons. This neural network has 1,856 variational parameters, which is more than two times fewer than the dimension of the Hilbert space for this case, which amounts to 4,845. We observed a rapid decline in the energy accuracy of NQS with increasing values of $\alpha$. Figure 1 illustrates the relationship between energy error (obtained by comparing with exact diagonalization) and $\alpha$. The energy error associated with NQS escalates by several orders of magnitude within the $0.20<\alpha<$ 0.2 range and experiences a further increase of an order of magnitude within the $0.30.2<\alpha<$ 0.3 range. Importantly, the observed decline in performance is virtually identical whether SR or SITE is used for NQS optimization. This uniformity implies that the underlying issue is unlikely to originate from the optimization algorithm itself.
@@
We also incorporated into Figure 1 the results obtained using the density matrix renormalization group (DMRG) method for comparison. For this we utilised TeNPy [58] Python package. We note that this is not the best result that could be achieved with DMRG as the virtual bond dimension of the Matrix Product State (MPS) was set to closely align the number of variational parameters (2,336) with that of our NQS. Notably, NQS shows significantly better performance at lower $\alpha$ values, becoming comparable to DMRG at higher values. While the performance of DMRG also deteriorates with increasing $\alpha$, this decline is considerably less pronounced than in the case of NQS. The employed virtual bond dimension of 8 is relatively small compared to the values typically used for DMRG. Therefore, we also included data with the bond dimension increased to 24. The performance improves significantly and surpasses NQS in the large magnetic flux regime. However, we note that in this case, the MPS contains 20,832 parameters, exceeding the Hilbert space dimension by about four times. Thus, even though the MPS has enough parameters to represent the ground state exactly, the DMRG algorithm fails to find such a solution, highlighting the well-known weakness of DMRG in approximating strongly entangled systems in more than one spatial dimension.
@@
We also observe a similar decrease in accuracy with increasing magnetic flux when directly maximizing the overlap between the NQS and the wave function obtained from ED (described in sec. 3.3.3). The overlap deviation from the target wave function vector increases by orders of magnitude in the regime of strong magnetic flux, compared to the scenario without a magnetic field. Remarkably, the specifics of the neural network architecture, as well as the size and density of the physical model, have minimal impact on this result. To validate this observation, we undertook an extensive series of experiments across a variety of neural network architectures, lattice dimensions, and particle counts. We explored every combination of the following parameters: the number of layers (2, 3, or 4); the number of neurons per hidden layer (32, 64, or 128); the number of lattice columns (4, 6, or 8); and the number of particles (4, 6, or 8). Additionally, we conducted analogous experiments using a CNN architecture, where we varied the number of channels (10, 20, or 44) instead of the number of neurons. The results from these experiments are illustrated in Fig. 1. Clearly, a relationship exists between the overlap deviation and the ratio of NQS parameters to the number of dimensions in the Hilbert space. The relationship is characterized by a significant dispersion of approximately one order of magnitude. This variability arises because different neural network configurations may vary in their optimality for this problem, and different physical system configurations may present varying levels of difficulty for approximation. Despite this large variation, a consistent overlap deviation gap of three to four orders of magnitude is observed between cases with $\alpha=0$ and $\alpha=0.3$.
@@
Figure 1: Left: Dependence of NQS and DMRG energy error on magnetic flux. The green curve represents DMRG with 8 virtual bonds (a similar number of variational parameters to NQS), while the orange curve corresponds to 24 virtual bonds (approximately 5 times more variational parameters than the Hilbert space dimension). Right: Overlap deviation plotted against the ratio of NQS parameters to the dimension of the Hilbert space. Each point represents a different run, varying in neural network architecture, lattice size, and particle number. Blue and red symbols indicate runs without magnetic fields and with strong magnetic fields, respectively. Circles and crosses denote runs using MLP and CNN neural network architectures, respectively.
Figure 2: Dependence of the required number of NQS parameters to achieve energy error $\Delta E_{\mathrm{req}}$ (10% of the energy gap between the ground and first excited states) as a function of magnetic flux. The blue, red, and green curves correspond to neural networks with depths of two, three, and four, respectively. The horizontal dashed line indicates the number of dimensions in the Hilbert space.
There is no general, well-established criterion for determining when the accuracy achieved by a numerical method is sufficient for the problem to be considered solved. However, it is worthwhile to investigate how the number of NQS parameters required to reach a certain accuracy depends on the value of the magnetic flux. For this purpose, we define the minimal required energy error as $\Delta E_{\mathrm{req}}=0.1\cdot(E_{\mathrm{exc}}-E_{\mathrm{gs}})$, where $E_{\mathrm{exc}}$ is the energy of the first excited state and $E_{\mathrm{gs}}$ is the energy of the ground state. Achieving this level of accuracy should ensure that the main features of the ground state are well approximated, and that the resulting state is clearly distinguishable from the excited states. Fig. 2 shows the number of NQS parameters, $n_{\mathrm{req}}$, required to achieve the specified accuracy as a function of magnetic flux $\alpha$. These experiments were conducted by fixing the depth of the MLP neural network to two, three, and four, while varying only the width. The exact values of $n_{\mathrm{req}}$ were obtained through linear interpolation. The number of required parameters increases by several orders of magnitude as $\alpha$ increases, and once again, the observed relationships are not sensitive to the network architecture (in this case, neural network depth). Interestingly, there are sudden jumps in $n_{\mathrm{req}}$ at $\alpha\approx 0.1$ and $\alpha\approx 0.2$, possibly indicating a transition after which the system becomes more difficult for NQS to approximate. Notably, when $\alpha>0.2$, the required number of parameters significantly exceeds the dimension of the Hilbert space.
@@ 4.2.4 Fractional quantum Hall effect
As it is well known [22], the noninteracting version of Hofstadter model is a paradigmatic model featuring the emergence of topological single-particle energy bands. In the presence of strong particle interactions and in the regime of a partially filled lowest Landau band, the Hofstadter Bose-Hubbard model has been conjectured to support the fractional Chern insulating (FCI) states, analogous to the Laughlin states in continuous systems [33, 34, 59, 60]. For bosons, the principal FCI state is ecpected at the filling factor of $\nu=1/2$ particles per flux quantum. Note that in finite systems with open boundaries, the number of plaquettes pierced by the flux is $(L_{x}-1)\times(L_{y}-1)$, thus the filling factor is defined [61] as $\nu=N/[\alpha(L_{x}-1)(L_{y}-1)]$, with $N$ denoting the number of particles. The formation of a fractionalized state on matter would certainly pose an additional (and unexplored) challenge for the NQS approach. However, we argue that in our case the performance drop cannot be ascribed to the stabilization of FCI states. In Figure 7 we plot the overlap deviation on the magnetic flux $\alpha$ (left panel) and the corresponding dependence on the filling factor $\nu$ (right panel) for a selection of lattice geometries. We see that the characteristic exponential growth of the overlap deviation is mainly associated with the regime of filling factors well above unity. The behavior of the overlap deviation below $\nu=1$ is rather erratic and the emerging peaks do not consistently point to a particular value of $\nu$, as one would expect looking for signals of FCI stabilization. All in all, we believe that in our experiments the considered regimes are not favourable for the emergence of FCI states, and they cannot be identified as a source of the performance gap.